A man's age is tree times his son's age ten years ago will be five times his son's age find both of their ages
Answer by Cromlix(4381)

Hi there,
A man's age is three times his son's age
M = 3S....(1)
ten years ago will be five times his son's age.
M - 10 = 5(S - 10)
M - 10 = 5S - 50......(20
Replace M in Eq (2) with 3S from Eq (1)
3S - 10 = 5S - 50
Collect like terms.
3S - 5S = -50 + 10
-2S = -40
S = -40/-2
S = 20
Man is 60 years old
Son is 20 years old.
Hope this helps :-)